OpenStudy (anonymous):

Is y=(x+5)/(x-3)(x+1) even, odd or neither?

OpenStudy (anonymous):

Is this \[y=\frac{x+5}{(x-3)(x+1)}\]

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

Yes, I mean't y=(x+3)/((x-3)(x+1))

OpenStudy (anonymous):

I also thought it was neither, because it has even and odd exponents, is my reasoning right or am I missing something?

OpenStudy (anonymous):

This is neither even nor odd

OpenStudy (anonymous):

\[y=\frac{x+5}{(x-3)(x+1)}\] is neither. you can check using the definition, or you can just pick a number for x. i pick 2 \[f(2)=\frac{2+5}{(2-3)(2+1)}=\frac{7}{-3}\] now i try with xm -2 and get \[\frac{-2+5}{(-2-5)(-2+1)}=\frac{3}{7}\] so not even or odd. if it was even i would get \[\frac{7}{-3}\] and if it were i should get \[\frac{7}{3}\]

OpenStudy (anonymous):

be careful about the exponents. \[\frac{x^2-4}{x^3}\] for example is odd, even though it has both even and odd exponents

OpenStudy (anonymous):

that is because \[x^2-4\] it even and \[x^3\] is odd and \[\frac{even}{odd}\] is odd.

OpenStudy (anonymous):

the real check is replacing x by -x and see if you get the same thing back. if so even. if you get the negative of the function back, it is odd. if you get neither, then neither even nor odd

OpenStudy (anonymous):

so for your original example you should check \[\frac{-x+5}{(-x-3)(-x+1)}\] and you will see that this is not \[\frac{x+5}{(x-3)(x+1)}\] nor it it \[-\frac{x+5}{(x-3)(x+1)}\]

OpenStudy (anonymous):

oh hello! did you stay up all night again?

myininaya (myininaya):

\[f(-x)=\frac{-x+5}{(-x-3)(-x+1)}=\frac{-x+5}{(-1)(x+3)(-1)(x-1)}=\frac{-x+5}{(x+3)(x-1)}\]

myininaya (myininaya):

yep

OpenStudy (anonymous):

beat you to it!

OpenStudy (anonymous):

i am off. see you later

myininaya (myininaya):

im a master :)

OpenStudy (anonymous):

\[\color{red}{\huge{\text{my myinninaya is a master!}}}\]

OpenStudy (anonymous):

except she is about to make a mistake. because \[-f(-x)\] is not what you wrote

myininaya (myininaya):

\[-f(x)=\frac{-(x+5)}{(x-3)(x+1)}\]

myininaya (myininaya):

if f(-x)=f(x) then it is even if f(-x)=-f(x) then it is odd if we have neither of the above ,it is neither so neither